OpenStudy (anonymous):

x-5=square root of x + 7

OpenStudy (saifoo.khan):

root x or root (x +7) ?

OpenStudy (anonymous):

\[x-5=\sqrt{x+7}\]

OpenStudy (anonymous):

\[x ^{2}-10x+25=x+7\] \[x ^{2}-11x+18=0\] \[(x-9)(x-2)\] \[x=9 , x=2\]